OpenStudy (anonymous):

Find the average rate of change from x = 7 to x = 14 for the function f(x) = 0.01(2)x and select the correct answer below. 23.22 35.84 128 163.84

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@mathmath333 @mathmate

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

i got c is that right

Nnesha (nnesha):

x is an exponent or is it 2 times x ?

OpenStudy (mathmate):

\(f(x)=0.01(2)^x\) \(f(7)=0.01(2)^7=1.28\) \(f(14)=0.01(2)^7=163.84\) Average rate of change of f(x) from 7 to 14 \(\Large =\frac{\Delta f(x)}{\Delta x}=\frac{f(14)-f(7)}{(14-7)}=?\)